OpenStudy (anonymous):

x - y = 2 0.2x + 0.4y = 1 The x-coordinate of the solution to the system shown is ___. 1 3 6

OpenStudy (anonymous):

let y=-2+x and plug it into the second equation to find x.

OpenStudy (anonymous):

Do you understand, or need me to show you?

OpenStudy (anonymous):

can u show me pls

OpenStudy (anonymous):

Ok. 0.2x + 0.4(-2+x)=1 0.2x - 0.8 + 0.4x =1 0.6x - 0.8= 1 0.6x = 1.8 x = 3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np